OpenStudy (anonymous):

I need help understanding the formula to calculate: If the top 10% of students are given A's, what is the lowest mark a student can get and still receive an A. Normal distribution, mean = 62 sd = 11. I can do the calculations but don't know how to set this up correctly. Thanks

OpenStudy (jack1):

call in the big guns... @jim_thompson5910 please help me @Jhannybean ...?

OpenStudy (jhannybean):

All the J people,eh?

OpenStudy (jhannybean):

Ew, standard deviation.

OpenStudy (jack1):

no, just u 2 because ur special ;P

OpenStudy (anonymous):

the instructor always uses something like P ( x < ----) P (Z < ----) or P (x > ---) P ( z > ____). Something to this effect...I don't know what way I am going, to the left, or right.

OpenStudy (jack1):

right, the top 10 %

jimthompson5910 (jim_thompson5910):

you want to find the value of k such that P(X > k) = 0.10

OpenStudy (anonymous):

My instructor has never used that formula...

jimthompson5910 (jim_thompson5910):

alternatively, you want to find the value of z such that P(Z > z) = 0.10 using mean = 0 and standard deviation = 1 once you get the critical value z, you would convert it to a raw score

OpenStudy (anonymous):

If I want the top 10%, do I need to find the z for 90th percentile and subtract from one?

jimthompson5910 (jim_thompson5910):

no just the z value for the 90th percentile

jimthompson5910 (jim_thompson5910):

the 90th percentile is the value that is above 90% of the distribution

OpenStudy (anonymous):

z for 90th percentile is 1.28

jimthompson5910 (jim_thompson5910):

ex: sometimes your grade or report card will say "you are in the 90th percentile", which would mean that your grade is higher than 90% of the student body

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

ie P(Z > 1.28) = 0.10 roughly

jimthompson5910 (jim_thompson5910):

now convert z = 1.28 to a raw score

jimthompson5910 (jim_thompson5910):

if you're not sure how to convert, then you just solve for x z = (x - mu)/sigma 1.28 = (x - 62)/11

OpenStudy (anonymous):

that is a similar formula he has had us use...this is what I came up with...1.28 (11) + 62 = 76.08 as the lowest possible score to achieve an A...but that doesn't sound reasonable

jimthompson5910 (jim_thompson5910):

76.08 is the correct answer

jimthompson5910 (jim_thompson5910):

btw, if you solve for x, you get 1.28 = (x - 62)/11 1.28*11 = x - 62 1.28*11 + 62= x x = 1.28*11 + 62

jimthompson5910 (jim_thompson5910):

which is what you did

OpenStudy (anonymous):

ok...thanks...you guys online write your formulas differently than he shows us in class...I am so confused..

OpenStudy (jack1):

quick query, what'd u use to get the sd value as 1.28 for 90%...?

OpenStudy (anonymous):

No...since I need to find my x value...I look for the 90th% ...the z value for .9000 = 1.28. then I solve for x by using 1.28 (11) + 62 ...(62 is mean) and 11 is sd...

OpenStudy (jack1):

yeah, but is there an equation to work out each of the percentiles? or is it a table you read off?

OpenStudy (anonymous):

sorry...I am NOT a math person and have no idea what I am doing....Yes, we have a Z table in our stats book... in fact...I am wondering if I'll get thru this class with a passing grade...when we are In class and do problems together, I get it...but later, it looks foreign...and I have so many pages of notes that I can't find what I'm looking for when I need it...

OpenStudy (jack1):

all good then, cheers beal and good luck thanks again for coming jim and Jhanny

jimthompson5910 (jim_thompson5910):

the formula you're using to convert is x = z*mu + sigma and that's just a rearranged version of what I originally wrote

OpenStudy (jack1):

... u sure it's not z* sigma + mu = x i had a look at the original above...

jimthompson5910 (jim_thompson5910):

oh sry, had them mixed up, yes it's x = z*sigma + mu

OpenStudy (jack1):

ah, cool

OpenStudy (anonymous):

Did I do it wrong then? He provided 2 formula's depending on what I need to find....Z = (x-mu)/sigma or X = mu + Z (sigma) ....but I still don't know how to set the problem up so I know what direction to go P ( X < or > of some number) = P (Z < or > of some number)....How can I know this and keep it straight?

jimthompson5910 (jim_thompson5910):

well you have a z chart, so use P(Z< ...) or P(Z >....)

jimthompson5910 (jim_thompson5910):

you want the top 10%, which means you want to find the value that will replace the question marks P(Z > ???) = 0.10

OpenStudy (anonymous):

So, although I want the top 10%, (Z > ??) I look at the 90th % to the left of Z. And if I wanted the least of some number, I would set it up as P (X < ?) P ( Z < ?)...right?

jimthompson5910 (jim_thompson5910):

yeah if you want the bottom 90%, then you want to find some number that replaces the question marks P(Z < ???) = 0.90